Part A Miller corporation has the following balance sheet (in E,000) ummary Balance Sheet ASSETS Cash Accounts reccivable Inventories Current Assets Net Fixed Assets Total Assets 30,000 60,000 60,000 150,000 150,000 300,000 LIABILITIES Accounts payable Accruals Short-term debt Current liabilities Long-term debt Preferred stock Common stock Retained earnings Total common equity Total liabilities and equity 30,000 30,000 15,000 75,000 90,000 15,000 30,000 90,000 120,000 300,000 The short-term debt is mainly IM Euro bank loans at 6%. These loans aim to finance receivables and inventories on a temporal basis. During the year its level is close to zero. The long-term debt consists of 7% coupon bonds, with a market price of 80 Euros. Face Value 100 Euros, and yield (YTM) 10%. The preferred stock has a par value of a 100 Euros, dividend yield of 12%, and trades in the market at 96 Euros. .The company has 12 million of common shares outstanding that trade at 20 Euros. The last dividend payment was 1.2 Euros, and dividends are expected to grow at an average growth rate of 8% The stocks beta is 1.5. The yield on government bonds is 7%, and the market risk premiun (Rm-Rf) is at 6%. Miller corporation is at the 40% tax bracket Estimate the WACC of Miller corporation using (a) book values (b) market values Note: all calculations should be provided

Part B 1. What sources of capital should be included in the estimation of the cost of capital (WACC) of a firm? (200 words) 2. In capital budgeting, should the costs be historical (embedded) costs or new (marginal) costs? (50 words)
Answer #1

Weighted Average Cost of Capital (“WACC”
is the ‘average of the cost’ of these sources of capital WACC is the minimum return that a company must earn on its existing capital to satisfy the respective stakeholder like shareholders, creditors, lenders etc. Mathematically, it is weighted average of the costs of each of the different types of capital which can be shown as the following equation:

Where:
E = Market value of equity
D = Market value of debt
P = Market value of preferred stock
V = E + D + P
Ke = Cost of equity
Kd = Cost of debt
Kp = Cost of preferred stock
t = tax rate

Calculation of WACC at Book Value:

Total Value of the capital(V)- Debt + Preferred Stock +Common Stock-90,000+15000+30000=135,000
Cost of equity = Rf + β × (Rm − Rf)- 7.5%+1.5(6%)=16.5%
Cost of debt = Interest rate × (1 – Corporate tax rate)
= 10 % × (1 – 40 %)=6%
Cost of preferred stock = 12 %
Therefore WACC at Book Value =90000/135000*16.5%+15000/135000*6%+30000/135000*12%=14.33%

Calculation of WACC at Market Value:

Total Value of the capital(V)- Debt + Preferred Stock +Common Stock-80,000+96,000+240000=416,000
Therefore WACC at Market Value-

WACC=240000/416000*16.5%+80000/416000*6%+96000/416000*12%=13.44%